Partners Networking to Advance South Island Dairying

Farm Walk Notes 18th March 2014

Posted by Rosie Petrowski on 19 March 2014

We have made some significant changes at LUDF over the last 2 weeks to enable the milking platform to stay within its historical Nitrogen load this season and to comply with the intent of the (ECAN) Land and Water Regional Plan.
 
We have so far dried off 74 MT’s and 9 culls. We have also put the small herd (234 earlier calving light BCS cows) onto OAD to ensure that they achieve their CS target at drying off and will achieve their CS targets at calving. Over the next week we will dry off a further 15 lower producing cows. Feeding the remaining cows 18kgDM/cow/day requires a growth rate of 60kg DM/ha/day; the stocking rate will now be 3.3 cows/ha. In essence LUDF is choosing to balance feed supply and demand by reducing demand, rather than increasing supply through importing more silage (with the subsequent additional N brought onto the platform).
 
The whole herd has gained 3kg live weight in the past week. The monitor group of 347 early MA cows have gained 2kg, and the younger cows [2011 born] have also gained 2kg live weight. Note: the make up of these groups has changed over the week as cows were dried off.
 
There are now 297 cows in the main herd and 234 in the small herd and 4 lame cows. This gives a total of 531 cows milking into the vat this morning.  There are 4 treatment cows. Average bulk milk SCC was 180.
 
Average daily milk solids production per cow (all cows milked into the vat) is 1.53kgMS/cow, 0.04 kg/ cow/day less than last week. Daily production per hectare was 5.08 kg MS/ha (vs 5.54kgMS/ha last week).
 
9 am average soil temperature was 16 degrees 2.5 degree warmer than last week. The farm had 8.2 mm rain over the week.
 
This weeks measured average pasture cover is 2482 kgDM/ha up 48kg DM/ha, from last week. Our estimated average daily pasture growth rate for the last week has been 76 kg DM/ha. This week we have a predicted surplus of 9 T DM. The round length has increased by 2 days over the week to 23 days. No silage has been fed this week. With what is looking like a cooler than average autumn setting in we will look to push the round out and hold it above 24 days this week. We have applied nitrogen [urea] to 37 ha of recently grazed pastures at 25kg N/ha.
 
The herd was body condition scored 10 days ago, average was 4.3 and 10% were below BCS 4.0 these are not as good as we would like. We had expected the average BCS to have started to trend upwards. We have already put low BCS, early calving cows in the small herd to assist with their body condition and will continue to monitor cows as ensure consistency of feeding towards BCS targets later in the season. We note that we have more low BCS cows so will be monitoring and acting on feeding thin cows intensively.